OpenStudy (anonymous):

Determine if the solution set for the system of equations shown is the empty set, contains one point or is infinite. x + y = 6 x - y = 0 a.one solution b.no solution c.infinite

OpenStudy (anonymous):

y = -x + 6 y = x two lines, different slopes, one intersection, one solution x = -x + 6 etc.